+ Half-life for First and Second Order Reactions 20 of 32 > A Review Constants Periodic Table ► View Available Hint(s) The ha<Topic_6_kinetics + Half-life for First and Second Order Reactions 20 of 32 > M Review l Constants Periodic Table For a firstThe half-life of a reaction, t1/2, is the time it takes for the reactant concentration [A] to decrease by half. For example, after one half-life the concentration falls from the initial concentration [A]0 to [A]0/2, after a second half-life to [A]0/4, after a third half-life to [A]0/8, and so on. on. For a first-order reaction, the half-life is constant. It depends only on the rate constant k and not on the reactant concentration. It is expressed as t1/2=0.693k For a second-order reaction, the half-life depends on the rate constant and the concentration of the reactant and so is expressed as t1/2=1k[A]0 Part A A certain first-order reaction (A→products) has a rate constant of 9.00×10−3 s−1 at 45 ∘C. How many minutes does it take for the concentration of the reactant, [A], to drop to 6.25% of the original concentration? Express your answer with the appropriate units. View Available Hint(s) nothing nothing Part B A certain second-order reaction (B→products) has a rate constant of 1.10×10−3 M−1⋅s−1 at 27 ∘C and an initial half-life of 218 s . What is the concentration of the reactant B after one half-life? Express your answer with the appropriate units. View Available Hint(s) nothing nothing Provide Feedback
Answer #1

PART-B * Half life of second order Reacties R !! Play {ts = [A].} O di k = 1.1x103 nt 5 th= 2185 (Ao - ? Just rearrange the eMelbod-1 ] Part A k=9x103s (A. = 100 (A) = 6.25 t= ? to *2.302 300 m - 1 *2.303 * log 10 x2.30 100 94103 = 2.303 xlog 16 9410